### v2

@@ Abstract
We study the online metric matching problem. There are $m$ servers and $n$ requests located in a metric space, where all servers are available upfront and requests arrive one at a time. Upon the arrival of a new request, it needs to be immediately and irrevocably matched to a free server, resulting in a cost of their distance. The objective is to minimize the total matching cost.
When servers are adversarial and requests are independently drawn from a known distribution, we reduce the problem to a more accessible setting where all servers and requests are independently drawn from the same distribution. Applying our reduction, for the Euclidean space $[0,1]^{d}$ with various choices of distributions, we achieve improved competitive ratios and nearly optimal regrets in both balanced and unbalanced markets. In particular, we give $O(1)$-competitive algorithms for $d\geq 3$ in both balanced and unbalanced markets with smooth distributions. Our algorithms improve on the $O((\log\log\log n)^{2})$ competitive ratio of Gupta et al. (ICALP’19) for balanced markets in various regimes, and provide the first positive results for unbalanced markets.
Moreover, when the servers and requests are adversarial and a prediction of request locations is provided, for balanced markets, we present a general framework for transforming an arbitrary algorithm that does not use predictions into an algorithm that leverages predictions. The competitive ratio of the resulting algorithm is a constant with a perfect prediction and degrades smoothly as the prediction accuracy deteriorates, while preserving the worst-case guarantee.
@@ 1 Introduction
The development of centralized matching markets has emerged as a crucial research area across operations research, economics, and computer science, driven by its diverse applications in kidney exchange programs, ride-sharing services, and online advertising platforms. Market designers in these domains face the fundamental challenge of making matching decisions under future uncertainty, with the primary objective of maximizing the total incurred value. The growing significance of these applications has sparked extensive research across multiple related areas. These include online bipartite matching [KVV90, Meh13, HTW24], online resource allocation [VB21, BKK23, BLM23, BBP24, Bra24, JMZ24], and dynamic matching [AAGK17, ALG20, WXY23, Gup24, KAG24, KYY25].
@@
A particularly compelling example of these applications can be found in the recent proliferation of ride-hailing platforms such as Didi, Uber, and Lyft. In these systems, the platform must make an immediate matching decision when a customer requests a ride - significantly delaying or rejecting the request is typically not an option. The primary cost metric in each match is the distance a driver needs to travel to reach their assigned customer. While a simple greedy approach matches each arriving customer to their nearest available driver, such myopic decisions can often lead to suboptimal outcomes. For instance, after matching a driver to their nearest customer, the next customer request might arrive in that driver’s original location, requiring a much longer pickup distance from the remaining available drivers. Therefore, the platform’s key challenge is to design matching policies that minimize the total pickup cost by considering both current matches and potential future request patterns, while maintaining the constraint of immediate matching for each request.
@@
Inspired by the special features of the market dynamics underlying ride-hailing platforms, in this paper, we focus on the online (bipartite) minimum cost matching problem. In this problem, $m$ nodes from one side of the market are called servers and are fully known upfront. On the other side of the market, which is often referred to as the request side, $n$ nodes arrive sequentially. We model the instantaneity of performing a match by assuming that each arriving request has to be immediately and irrevocably assigned to an available server. Due to the importance of application to ride-hailing, we assume that all servers and requests are located in the same metric space, and the cost incurred by matching a server and a request is measured by the distance between them. After realizing a match, both the matched server and request are removed. The objective of a central market planer is to minimize the total matching cost. We will consider both balanced and unbalanced markets, where balanced markets contain an equal number of servers and requests, and unbalanced markets possess at most a linear amount of excess servers. We will adopt competitive ratio (resp. regret) as our benchmark, which is defined as the ratio (resp. difference) between the expected average cost of matching each request incurred by the algorithm and that incurred by the optimal offline solution.
@@
A majority of literature on online metric matching focuses on the adversarial model [PS23, Rag18, MNP06], where the locations of all servers and requests are adversarially chosen, or the weaker random-order model [Rag16], which assumes that the arrival order of adversarially chosen requests is randomly permuted. Both models, albeit being fairly general, suffer from super-constant competitive ratio lower bounds [Rag16, PS23]. Recent lines of work bypass these barriers by modeling more practical and tractable scenarios in which algorithms are equipped with additional information regarding future uncertainty. Specifically, [GGPW19, BFP23] assume that requests are randomly drawn from a known distribution, and [ACE+23b] consider the case where algorithms are provided with predictions of request locations. They achieve improved or even nearly optimal results under various choices of modeling assumptions.
@@ 1.1 Main Contribution
As our main contribution, we propose a generic approach to designing algorithms for the online metric matching problem that leverages information about future uncertainty - either through probability distributions or predictions of request locations. These algorithms result from enhancing an arbitrary algorithm ${\mathcal{A}}$ that does not use additional information. A key advantage of our approach is that this enhancement only invokes ${\mathcal{A}}$ in a black-box manner without imposing technical assumptions on ${\mathcal{A}}$.
More specifically, given predictions of request locations, for an arbitrary algorithm ${\mathcal{A}}$ that does not use predictions, we can convert ${\mathcal{A}}$ into an enhanced algorithm ${\mathcal{A}}^{\prime}$ that bases its matching decisions on both the predictions and ${\mathcal{A}}$’s strategy. The performance of ${\mathcal{A}}^{\prime}$ depends jointly on the prediction accuracy and ${\mathcal{A}}$’s performance (Theorem 1). When given distributional information instead, we generate predictions by sampling from the distribution, allowing us to apply the same approach.
We then demonstrate the effectiveness of our framework through several significant results in the stochastic and prediction settings respectively, following by our additional results for the adversarial setting.
Stochastic Setting.
When servers are adversarially chosen and requests are independently drawn from a known distribution $\mathbb{D}$, we show that the aforementioned approach results in a reduction to the setting with stochastic servers and requests. Specifically, if there is an algorithm ${\mathcal{A}}$ for balanced markets with servers and requests all independently drawn from $\mathbb{D}$, we can use ${\mathcal{A}}$ in a black-box manner to derive another algorithm ${\mathcal{A}}^{\prime}$ for unbalanced markets with servers adversarially chosen and requests independently drawn from $\mathbb{D}$ (Theorem 2). Moreover, if the markets are balanced, which can be viewed as a special case of unbalanced markets, ${\mathcal{A}}^{\prime}$ admits the same competitive ratio as ${\mathcal{A}}$ up to a constant factor. Next, we describe our results in more concrete settings obtained by further utilizing the recent progress in the setting with stochastic servers and requests [Kan22, CKKZ23].
@@
Table 1: Competitive ratios for the Euclidean space $[0,1]^{d}$ with adversarial servers and stochastic requests. All previous state-of-the-art results are given in black fonts, where we use $\omega(1)$ to denote $O((\log\log\log n)^{2})$, and all new results achieved in this paper are marked in red. In particular, for balanced markets with general distributions, the competitive ratios of $O(1)$ for $d=1$ and of $O((\log\log\log n)^{2})$ for $d\geq 2$ are obtained by [GGPW19], which imply the bounds on their left.
@@
We first introduce the competitive ratio results for the Euclidean space $[0,1]^{d}$ achieved by combining our reduction with algorithms of prior work, which are summarized in Table 1. Prior to our work, the best algorithm in balanced markets with general metrics and distributions achieves a competitive ratio of $O((\log\log\log n)^{2})$, and the same algorithm is known to be $O(1)$-competitive for tree metrics [GGPW19]. We start by showing that combining our reduction with the Hierarchical Greedy algorithm of [Kan22] leads to an $O(1)$-competitive algorithm in balanced markets for all $d\geq 1$ with adversarial servers and uniformly distributed requests (Theorem 4).
@@
Next, we describe our results in balanced markets with servers adversarially chosen and requests independently drawn from a smooth distribution, where a distribution is smooth if it admits a density function upper bounded by a constant. For the algorithm ${\mathcal{A}}$ required by the reduction, we adopt the Simulate-Optimize-Assign-Repeat (SOAR) algorithm of [CKKZ23], which results in an $O(1)$-competitive algorithm for $d\geq 3$ (Theorem 5).
@@
Finally, we present our results in unbalanced markets with adversarial servers and stochastic requests. When requests are independently drawn from a smooth distribution, by applying the same approach as for balanced markets, the resulting algorithm admits a competitive ratio of $O(\sqrt{n})$ for $d=1$, $O(\sqrt{\log n})$ for $d=2$, and $O(1)$ for $d\geq 3$ (Theorem 6). To improve the undesirable competitive ratio for $d=1$, we show that the $O(\log n)$ competitive ratio of the algorithm proposed by [GL12], which holds in balanced markets when all servers and requests are adversarially located in a doubling metric [^1] A metric space is said to be doubling if a doubling constant $M>0$ exists such that for all $r>0$, every ball of radius $r$ can be covered by at most $M$ balls of radius $r/2$. In particular, the Euclidean space $[0,1]^{d}$ is a doubling metric for constant $d$., is preserved in unbalanced markets (Corollary 3). These constitute the first known results in unbalanced markets for adversarial servers and stochastic requests.
@@
We further describe our additional results with respect to regret analysis, which are for the Euclidean space $[0,1]^{d}$. Recall that in balanced markets with servers and requests following different and arbitrary distributions, [CKKZ23] show that the SOAR algorithm admits a regret of $O(n^{-\frac{1}{2}})$ for $d=1$, $O(n^{-\frac{1}{2}}\log n)$ for $d=2$, and $O(n^{-\frac{1}{d}})$ for $d\geq 3$, which is optimal for $d\neq 2$ and optimal up to a logarithmic factor for $d=2$. By combining our reduction with the SOAR algorithm, we show that the same regrets are achievable even in unbalanced markets with servers adversarially chosen and requests independently drawn from an arbitrary distribution (Theorem 7). Since balanced markets form a special case of unbalanced markets, and our setting with adversarial servers is more general, the lower bounds by [CKKZ23] continue to hold in our case, implying our regret guarantees to be nearly optimal.
@@
In the prediction setting, algorithms receive a set of $n$ points as a prediction at the beginning, which predicts the unknown request locations. We will focus on balanced markets with servers and requests being adversarially chosen. In particular, the prediction is not guaranteed to be correct with its error oblivious to algorithms, and algorithms are expected to be (1) consistent: the performance of algorithms is nearly optimal with a perfect prediction, and (2) robust: when the prediction admits a large error, the performance of algorithms is comparable to that of the best online algorithm that does not use predictions. In our case, the prediction error $\eta$ is defined as the minimum cost of all perfect matchings between the set of predicted request locations and the set of actual request locations.
Given an $\alpha$-competitive algorithm ${\mathcal{A}}$ that does not use predictions, we apply our generic approach mentioned at the beginning of this subsection to derive an algorithm ${\mathcal{A}}^{\prime}$ that leverages predictions and achieves a competitive ratio of $O(\min\{\alpha,1+\frac{\eta(\alpha+1)}{\mathrm{OPT}}\})$, where $\mathrm{OPT}$ denotes the optimal offline cost (Theorem 8). To see the consistency of ${\mathcal{A}}^{\prime}$, its competitive ratio converges to a constant when the prediction error $\eta$ approaches 00. Moreover, the competitive ratio of ${\mathcal{A}}^{\prime}$ degrades smoothly as the prediction error increases, and the competitive ratio guarantee of $\alpha$ provided by ${\mathcal{A}}$ is maintained up to a constant factor regardless of the prediction error, indicating the robustness of ${\mathcal{A}}^{\prime}$. Finally, we apply the above result to several specific settings (Corollary 1).
@@
Now, we discuss our results for the setting in which servers and requests are adversarially chosen with no predictions provided. As alluded to previously, we generalize the $O(\log n)$ competitive ratio of the algorithm by [GL12] for doubling metrics from balanced markets to unbalanced markets. This is achieved based on the observation that every greedy algorithm with a tie-breaking rule satisfying certain properties admits the same competitive ratio in balanced and unbalanced markets (Theorem 9). As a by-product, we similarly show that the $O(\log^{3}n)$ competitive ratio of the algorithm by [MNP06] in balanced markets with servers and requests adversarially located in an arbitrary metric is retained in unbalanced markets (Corollary 2). We remark here that all prior results for adversarial servers and requests are established in balanced markets, and it is not immediate whether they can be carried over to unbalanced markets.

### v3

@@ Abstract
We study the online metric matching problem. There are $m$ servers and $n$ requests located in a metric space, where all servers are available upfront and requests arrive one at a time. Upon the arrival of a new request, it needs to be immediately and irrevocably matched to an available server, resulting in a cost of their distance. The objective is to minimize the total matching cost.
When servers are adversarial and requests are independently drawn from a known distribution, we reduce the problem to a more tractable setting where servers and requests are all independently drawn from the same distribution. Applying our reduction, for $0,1]^{d}$ with various choices of distributions, we achieve improved competitive ratios and nearly optimal regret in both balanced and unbalanced markets. In particular, we give $O(1)$-competitive algorithms for $d\geq 3$ in both balanced and unbalanced markets with smooth distributions. Our algorithms improve on the $O((\log\log\log n)^{2})$ competitive ratio of Gupta et al. (ICALP’19) for balanced markets in various regimes, and provide the first positive results for unbalanced markets. Moreover, when servers and requests are all adversarial, and a prediction of request locations is provided, we present a general framework for transforming an arbitrary algorithm that does not use predictions into an algorithm that leverages predictions. The transformation applies the given algorithm in a black-box manner, and the performance of the resulting algorithm degrades smoothly as the prediction accuracy deteriorates while preserving the worst-case guarantee.
@@ 1 Introduction
The development of centralized matching markets has emerged as a crucial research area across operational research, economics, and computer science, driven by its diverse applications in kidney exchange programs, ride-sharing services, and online advertising platforms. Market designers in these domains face the fundamental challenge of making matching decisions under future uncertainty, with the primary objective of maximizing the total incurred value. The growing significance of these applications has sparked extensive research across multiple related areas. These include online bipartite matching [KVV90, Meh13, HTW24], online resource allocation [VB21, BKK23, BLM23, BBP24, Bra24, JMZ24], and dynamic matching [AAGK17, ALG20, WXY23, Gup24, KQYY25].
@@
A particularly compelling example of these applications can be found in the recent proliferation of ride-hailing platforms such as Didi, Uber, and Lyft. In these systems, the platform must make an immediate matching decision when a customer requests a ride—significantly delaying or rejecting the request is typically not an option. The primary cost metric in each match is the distance a driver needs to travel to reach their assigned customer. While a simple greedy approach matches each arriving customer to their nearest available driver, such myopic decisions can often lead to suboptimal outcomes. For instance, after matching a driver to their nearest customer, the next customer request might arrive in that driver’s original location, requiring a much longer pickup distance from the remaining available drivers. Therefore, the platform’s key challenge is to design matching policies that minimize the total pickup costs by considering both current matches and potential future request patterns, while maintaining the constraint of immediate matching for each request.
@@
Inspired by the special features of the market dynamics underlying ride-hailing platforms, in this paper, we focus on the online (bipartite) minimum cost matching problem. In this problem, $m$ nodes from one side of the market are called servers and are fully known upfront. On the other side of the market, which is often referred to as the request side, $n$ nodes arrive sequentially. We model the instantaneity of performing a match by assuming that each arriving request has to be immediately and irrevocably assigned to an available server. Due to the importance of application to ride-hailing, we assume that all servers and requests are located in the same metric space, and the cost incurred by a match is measured by the distance between two matched nodes. After realizing a match, both the matched server and request are removed. The objective of a central market planer is to minimize the total matching cost.
Beyond ride-hailing, online metric matching also finds applications in service platforms, online gaming, and kidney exchanges. In these settings, participants’ features are encoded as high-dimensional vectors, where the distance measures incompatibility—for instance, skill levels in gaming matchmaking, medical compatibility factors in kidney exchanges, or service requirements in platform matching.
We consider both balanced and unbalanced markets, where balanced markets contain an equal number of servers and requests, and unbalanced markets possess more servers than requests. Note that balanced markets constitute a special case of unbalanced markets. We will adopt competitive ratio (resp. regret) as our performance measure, which is defined as the ratio (resp. difference) between the expected average cost of matching each request incurred by the algorithm and that incurred by the optimal offline solution.
@@
A majority of literature on online metric matching focuses on the adversarial model [MNP06, Rag18, PS23], where the locations of all servers and requests are adversarially chosen, or the random-order model [Rag16], which assumes that the arrival order of adversarially chosen requests is randomly permuted. Both models, albeit being fairly general, suffer from super-constant competitive ratio lower bounds [MNP06, Rag16, PS23]. Recent lines of work bypass these barriers by modeling more practical and tractable scenarios in which algorithms are equipped with additional information on future uncertainty. Specifically, [GGPW19, BFP23] assume that requests are independently drawn from a known distribution, and [ACE+23b] consider the case where algorithms are provided with predictions of the behaviors of the optimal offline algorithm. They achieve improved or even nearly optimal results under these weaker models.
@@ 1.1 Main Contribution
As our main result, we propose a generic framework to designing algorithms for online metric matching that leverages information on future uncertainty provided in the form of predicted request locations. Our framework operates by enhancing an arbitrary algorithm $\mathcal{A}$} that does not use additional information. As a key feature of our framework, the enhancement only invokes $\mathcal{A}$} in a black-box manner without imposing technical assumptions on $\mathcal{A}$}.
More specifically, given a prediction of request locations, for an arbitrary algorithm $\mathcal{A}$} that does not use predictions, we can convert $\mathcal{A}$} into an enhanced algorithm $\mathcal{A}$$^{\prime}$ that bases its matching decisions on both the prediction and $\mathcal{A}$}’s strategy, and the performance of $\mathcal{A}$$^{\prime}$ depends jointly on the prediction accuracy and $\mathcal{A}$}’s performance (Theorem 1).
We then demonstrate the effectiveness of our framework by applying it to the semi-stochastic and prediction settings, following by our additional results for the adversarial setting.
Semi-Stochastic Setting.
For the semi-stochastic setting where servers are adversarially chosen and requests are independently drawn from a known distribution $\mathbb{D}$, by sampling from $\mathbb{D}$} to generate a prediction of request locations, we apply our generic framework to reduce the semi-stochastic setting to the stochastic setting where all servers and requests are independently drawn from $\mathbb{D}$. Specifically, given an algorithm $\mathcal{A}$} for the stochastic setting and balanced markets, we can use $\mathcal{A}$} in a black-box manner to derive an algorithm $\mathcal{A}$$^{\prime}$ for the semi-stochastic setting and unbalanced markets (Theorem 2). Moreover, if the markets for the semi-stochastic setting are balanced, $\mathcal{A}$$^{\prime}$ inherits the competitive ratio of $\mathcal{A}$} up to a constant factor. Next, we describe our concrete results in the semi-stochastic setting obtained by combining our reduction with the recent progress in the stochastic setting [Kan22, CKKZ23].
@@
Table 1: Competitive ratios for the Euclidean metric $0,1]^{d}$ in the semi-stochastic setting. All previous state-of-the-art results are given in black fonts, where we use $\omega(1)$ to denote $O((\log\log\log n)^{2})$, and all new results achieved in this paper are marked in red. In particular, for balanced markets with general distributions, the competitive ratios of $O(1)$ for $d=1$ and of $O((\log\log\log n)^{2})$ for $d\geq 2$ are achieved by [GGPW19], which imply the bounds on their left.
@@
We first introduce the competitive ratio results for the Euclidean metric $0,1]^{d}$, which are summarized in Table 1. Prior to our work, the best competitive ratio achieved for balanced markets with general metrics and distributions is $O((\log\log\log n)^{2})$, and the same algorithm is $O(1)$-competitive for tree metrics [GGPW19]. We start by showing that combining our reduction with the algorithm of [Kan22] leads to an $O(1)$-competitive algorithm for all $d\geq 1$ in balanced markets and the semi-stochastic setting with the uniform distribution (Theorem 3).
@@
Next, we describe our results in balanced markets and the semi-stochastic setting with smooth distributions, where a distribution is smooth if it admits a density function upper bounded by a constant. For the algorithm $\mathcal{A}$} required by the reduction, we adopt the Simulate-Optimize-Assign-Repeat (SOAR) algorithm of [CKKZ23], which results in an $O(1)$-competitive algorithm for $d\geq 3$ (Theorem 5).
@@
Finally, we present our results for the semi-stochastic setting and unbalanced markets, for which we assume that servers outnumber requests by at most a constant factor. For smooth distributions, by applying the same approach as for balanced markets, the resulting algorithm admits a competitive ratio of $O(\sqrt{n})$ for $d=1$, $O(\sqrt{\log n})$ for $d=2$, and $O(1)$ for $d\geq 3$ (Theorem 6). To improve the undesirable bound for $d=1$, we show that the $O(\log n)$ competitive ratio of the algorithm proposed by [GL12], which holds in balanced markets when servers and requests are all adversarially located in a doubling metric [^1] A metric space is doubling if there exists a constant $M>0$ such that for all $r>0$, every ball of radius $r$ can be covered by at most $M$ balls of radius $r/2$. In particular, the Euclidean metric $0,1]^{d}$ is doubling for constant $d$., is preserved in unbalanced markets (Corollary 3). These constitute the first known results for the semi-stochastic setting and unbalanced markets.
Stoch. and Balanced Semi-Stoch. and Unbalanced
Lower Bound Upper Bound Lower Bound Upper Bound
$d=1$ $\Omega(n^{-\frac{1}{2}})$ $O(n^{-\frac{1}{2}})$ $\Omega(n^{-\frac{1}{2}})$ $O(n^{-\frac{1}{2}})$
$d=2$ $\Omega(n^{-\frac{1}{2}})$ $O(n^{-\frac{1}{2}}\sqrt{\log n})$ $\Omega(n^{-\frac{1}{2}})$ $O(n^{-\frac{1}{2}}\sqrt{\log n})$
$d\geq 3$ $\Omega(n^{-\frac{1}{d}})$ $O(n^{-\frac{1}{d}})$ $\Omega(n^{-\frac{1}{d}})$ $O(n^{-\frac{1}{d}})$
Table 2: Regret for the Euclidean metric $0,1]^{d}$ with general distributions. All previous results are given in black fonts, and all new results achieved in this paper are marked in red. In particular, the results for the stochastic setting and balanced markets are achieved by [CKKZ23], which imply the lower bounds for the semi-stochastic setting and unbalanced markets.
@@
We further describe our results on regret analysis for the Euclidean metric $0,1]^{d}$, which are summarized in Table 2. Recall that for balanced markets and the stochastic setting with general distributions, [CKKZ23] show that the SOAR algorithm admits regret of $O(n^{-\frac{1}{2}})$ for $d=1$, $O(n^{-\frac{1}{2}}\sqrt{\log n})$ for $d=2$, and $O(n^{-\frac{1}{d}})$ for $d\geq 3$, which is optimal for $d\neq 2$ and optimal up to a logarithmic factor for $d=2$. By combining our reduction with the SOAR algorithm, we show that the same regret is attainable even for unbalanced markets and the semi-stochastic setting with general distributions (Theorem 7). Since the case with balanced markets and the stochastic setting constitutes a special case of unbalanced markets and the semi-stochastic setting, the lower bounds by [CKKZ23] imply our regret guarantees to be nearly optimal.
@@
In the prediction setting, the algorithm receives a set of $n$ points as a prediction in advance, which predicts the unknown request locations. We consider balanced markets and the adversarial setting, where servers and requests are all adversarially chosen. In particular, the prediction is not guaranteed to be accurate, and its error is oblivious to the algorithm. The algorithm is expected to be (1) consistent: its performance is nearly optimal with perfectly accurate predictions, and (2) robust: with inaccurate predictions, its performance is comparable to the best online algorithm that does not use predictions. In our case, the prediction error $\eta$ is defined as the minimum cost of all perfect matchings between the prediction and the realized request locations.
Given an $\alpha$-competitive algorithm $\mathcal{A}$} that does not use predictions, we apply our generic framework to derive an algorithm $\mathcal{A}$$^{\prime}$ that leverages the prediction and achieves a competitive ratio of $O(\min\{\alpha,1+\frac{\eta(\alpha+1)}{\mathrm{OPT}}\})$, where $\mathrm{OPT}$ denotes the optimal offline cost (Theorem 8). To see the consistency of $\mathcal{A}$$^{\prime}$, its competitive ratio converges to a constant when the prediction error $\eta$ approaches 0. Moreover, under a large prediction error, we recover the competitive ratio $\alpha$ provided by $\mathcal{A}$} up to a constant factor, fulfilling the robustness requirement. Besides, the performance of $\mathcal{A}$$^{\prime}$ degrades smoothly as the prediction error increases, which is usually referred to as the smoothness property. Finally, we apply the above result to several specific metrics (Corollary 1).
To highlight the practical relevance of our prediction model, in reality, ride-hailing platforms may not predict exact future request locations but rather use historical data to estimate demand distributions across regions. Our prediction model with point locations serves as a useful abstraction that captures various practical scenarios: (1) predictions could be sampled from estimated distributions, providing robustness when these estimates are imperfect—unlike the semi-stochastic setting which assumes perfect distributional knowledge; (2) predictions could represent anticipated requests based on scheduled events or historical patterns; (3) the prediction error $\eta$ naturally quantifies uncertainty in any forecasting method.
Our framework’s strength lies in its generality—it provides performance guarantees regardless of how predictions are generated, degrading gracefully from near-optimal performance with accurate predictions to worst-case guarantees with poor predictions. This robustness is particularly valuable when distributional assumptions may not hold perfectly in practice.
@@
We turn to our results for the adversarial setting without predictions. As alluded to previously, we generalize the $O(\log n)$ competitive ratio of the algorithm by [GL12] for doubling metrics from balanced markets to unbalanced markets. This is achieved based on the more general result that every greedy algorithm with a tie-breaking rule satisfying certain properties admits the same competitive ratio in balanced and unbalanced markets (Theorem 9). As a by-product, we also show that the $O(\log^{3}n)$-competitive algorithm by [MNP06] for balanced markets with general metrics retains its competitive ratio in unbalanced markets (Corollary 2). We remark here that all prior results in the adversarial setting are established in balanced markets, and it is not immediate whether they can be carried over to unbalanced markets.